OpenStudy (anonymous):

(3 - 13x - 7x2) - (5x2 + 12x - 10) =

OpenStudy (anonymous):

another one? ok here we go, ill walk your through it

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

First you need to get rid of the parentheses. The first set of parentheses is unnecessary, so just drop them. To get rid of the second set of parentheses, distribute the negative outside by all terms inside. The negative outside is like -1, so multiply every term inside the second set of parentheses by -1.

OpenStudy (anonymous):

so distribute the negative into the problem after it

OpenStudy (goldphenoix):

For the second set of parenthesis, you would just need to change it to the opposoite operation. So ex: Opposite of addition is subtraction Opposite of subtraction is addition.

OpenStudy (anonymous):

(3 - 13x - 7x2) (-5x2 - 12x + 10) would be your new equation

OpenStudy (goldphenoix):

Not quite. Take out the parenthesis, then it would be correct.

OpenStudy (anonymous):

yeah sorry my bad then you would just combine like terms

OpenStudy (goldphenoix):

So it would look like: \[\large \large 3 - 13x - 7x^2 -5x^2 - 12x + 10\]

OpenStudy (anonymous):

Gold i need help with my newest question care to help after this

OpenStudy (goldphenoix):

I'll try.

OpenStudy (anonymous):

12x^2 - 25x + 13 would be your answer

OpenStudy (goldphenoix):

Not quite. -7x^2 + -5x^2 = -12x^2

OpenStudy (anonymous):

-7x^ - 5x^ = -12x^2

OpenStudy (mathstudent55):

\((3 - 13x - 7x^2) - (5x^2 + 12x - 10) =\) The first set of parentheses is just dropped: \(= 3 - 13x - 7x^2 - (5x^2 + 12x - 10)\) Every term inside the second set of parentheses is multiplied by -1: \(= 3 - 13x - 7x^2 - 5x^2 -12x +10\) Combine like terms: \(= 3 + 10 -13x - 12x - 7x^2 - 5x^2\) \(= 13 - 25x - 12x^2\)

OpenStudy (anonymous):

@mathstudent55 just remember that the highest variable goes first so just write the equation you wrote backwards